Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purchase is completed on January 1, 2017, at a total cash price of $800,000 for a building, land, land improvements, and four vehicles. The estimated market values of the assets are building, $482,500; land, $308,800; land improvements, $38,600; and four vehicles, $135,100. The company’s fiscal year ends on December 31.

Required:

1-a. Prepare a table to allocate the lump-sum purchase price to the separate assets purchased.

1-b. Prepare the journal entry to record the purchase.

2. Compute the depreciation expense for year 2017 on the building using the straight-line method, assuming a 15-year life and a $31,000 salvage value.

3. Compute the depreciation expense for year 2017 on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1
1a
Allocation of total cost Appraised Value Percent of Total
Appraised Value
x Total cost of
Acquisition
Apportioned Cost
Building 482,500 50% x 800,000 400,000
Land 308,800 32% x 800,000 256,000
Land improvements 38,600 4% x 800,000 32,000
Vehicles 135,100 14% x 800,000 112,000
Total 965,000 100% 800,000
b
Date General Journal Debit Credit
1-Jan Building 400,000
Land 256,000
Land improvements 32,000
Vehicles 112,000
Cash 800,000
2
Depreciation expense on building 24600 =(400000-31000)/15
3
Depreciation rate 40% =1/5*2
Depreciation expense on
land improvements
12800 =32000*40%